Bonferroni Company sells 1,000 Halloween costumes to a wholesaler for $20 per costume. The contract provides the wholesaler the right to return all unsold costumes after Halloween. Based on its experience in this industry, Bonferroni can reliably provide the following range of returns for this contract.

1. What is the estimated transaction price if Bonferroni uses the expected value method? What is the estimated transaction price if Bonferroni uses the most likely amount method?

Number of Costumes Returned - Probability

50 - 40%

75 - 45%

100 - 15%
Answer #1

QUESTION

Bonferroni Company sells 1,000 Halloween costumes to a wholesaler for $20 per costume. The contract provides the wholesaler the right to return all unsold costumes after Halloween. Based on its experience in this industry, Bonferroni can reliably provide the following range of returns for this contract.

1. What is the estimated transaction price if Bonferroni uses the expected value method? What is the estimated transaction price if Bonferroni uses the most likely amount method?

Number of Costumes Returned - Probability

50 - 40%

75 - 45%

100 - 15%

  SOLUTION

  1. The estimated transaction price if Bonferroni uses the expected value method:-

Calculation of Expected Number of Costume returns :-

No. of Returns (X)

Probability (P)

(P)*(X)

50

40%

20.00

75

45%

33.75

100

15%

15.00

ΣPX

68.75

So, Number of Costumes Sold = 1,000

Number of Costumes Returned = 68.75 = 69 (Rounded off as no. of Returns cant be in a fraction)

Net number of costumes sold = 1000-69 = 931

Sales Price = $ 20.00

So, Estimated transaction price = Number of Costumes Sold * Sales Price

= 931 * $ 20.00

= $ 18,620.00

NOTE:-

Had we not ignored the Fraction in number of Returns,

Net number of costumes sold would have been = 1000 - 68.75 = 931.25

And, Estimated transaction price = 931.25 * $ 20 = $ 18,625.00

  1. The estimated transaction price if Bonferroni uses the most likely amount method?

Since most likely number of Returns is “75” i.e. with highest probability of 45%,

Net number of Costumes Sold = 1000-75 = 925

Estimated transaction price = Number of Costumes Sold * Sales Price

= 925 * $ 20.00

= $ 18,500.00